OpenStudy (anonymous):

what is 5% of 120???

OpenStudy (lgbasallote):

hint \[5 \text % \; \text{of}\; 120 \implies 0.05 \times 120\] does that help?

OpenStudy (anonymous):

divide 5 by 120

OpenStudy (anonymous):

so it is 6??

OpenStudy (anonymous):

then move the decimal to the twice

OpenStudy (anonymous):

*to the left

OpenStudy (lgbasallote):

yes 6

OpenStudy (anonymous):

thank you!!!!

OpenStudy (lgbasallote):

dividing 5 by 120 would be "what percent of 120 is 5" right?

OpenStudy (anonymous):

what is 60% of 1000

OpenStudy (lgbasallote):

what do you think?

OpenStudy (anonymous):

idk

OpenStudy (lgbasallote):

do the same as i did before

OpenStudy (anonymous):

6

OpenStudy (lgbasallote):

divide 60 by 100 then multiply to 1000

OpenStudy (lgbasallote):

\[\implies \frac{60}{100} \times 1000\]

OpenStudy (anonymous):

6000

OpenStudy (anonymous):

600

OpenStudy (lgbasallote):

is that your final guess?

OpenStudy (anonymous):

yes 600

OpenStudy (lgbasallote):

right

OpenStudy (anonymous):

yes!!!!!

OpenStudy (anonymous):

Sales tax is calculated as a percent of the sales price. If sales tax is 6%, what is the sales tax on clothing that costs $180?

OpenStudy (anonymous):

what the answer???? or hint???

OpenStudy (lgbasallote):

sorry. i dont know much about taxes. i usually dont evade them.

OpenStudy (anonymous):

awwww ok

OpenStudy (anonymous):

wait

OpenStudy (lgbasallote):

however this seems like it's just 6% of180

OpenStudy (anonymous):

50% of 700

OpenStudy (lgbasallote):

what do you think is 50% of 700?

OpenStudy (anonymous):

500

OpenStudy (lgbasallote):

how did you get that again?

OpenStudy (anonymous):

650????

OpenStudy (lgbasallote):

are we talking about 50% of 700?

OpenStudy (anonymous):

yes!!

OpenStudy (lgbasallote):

so what did you do to get those answers?

OpenStudy (anonymous):

idk i guessed

OpenStudy (lgbasallote):

to turn percent into fraction divide by 100. for example 70% of 700 is \[\frac{70}{100} \times 700\] so what is 50% of 700 then?

OpenStudy (anonymous):

490

OpenStudy (lgbasallote):

that is 70% of 700......

OpenStudy (lgbasallote):

if 70% is \[\frac{70}{100}\] then what is 50%?

OpenStudy (anonymous):

350